evaluate f(-3) for the function f(x) = 5x2(squared) - 2x - 4
Answer by edjones(8007)

f(x) = 5x^2 - 2x - 4
f(3)=5(-3)^2-2(-3)-4
f(3)=5*9+6-4
f(3)=45+6-4=47
Ed